# A User-Friendly Framework for Generating Model-Preferred Prompts in Text-to-Image Synthesis

### 20 Feb 2024 v1

@@ Introduction
Generative foundation models, including language models and text-to-image models, can be prompted to follow user instructions. Recent advancements in text-to-image synthesis such as Stable Diffusion (SD) (Rombach et al. 2022) and Midjourney (Holz 2023) have facilitated the generation of high-fidelity images based on text prompts. Concurrently, recent studies have revealed that prompt design plays a crucial role in determining the quality of the generated images (Ko et al. 2023; Liu, Qiao, and Chilton 2022). Adjusting the prompt to better reflect the user’s intentions can lead to superior results. This issue is particularly pronounced in text-to-image models, as the capacity of their text encoders is relatively limited, such as CLIP text encoder (Radford et al. 2021) in Stable Diffusion. Empirical observations have also shown that common user inputs are often insufficient to produce aesthetically pleasing images using current models (Hao et al. 2023).
@@
In the specific field of text-to-image generative models, prompt engineering research is nascent, which aims to use carefully selected and combined sentences to achieve a specific visual style in synthesized images (Oppenlaender 2023). The input texts, known as the prompts, direct the model to generate specific images. Manual prompt engineering is a natural method for optimizing prompts. Pavlichenko et al. (2023); Oppenlaender et al. (2023); Liu et al. (2022) explore how to find model-preferred prompts manually. While manual prompt engineering can lead to significant progress, the process of designing prompts requires time and experience, and may not always yield optimal results. Therefore, various methods have focused on automatically searching for prompts through mining (Jiang et al. 2020), parsing (Haviv, Berant, and Globerson 2021), text generation (Hao et al. 2023) and LLMs (Zhu et al. 2023; Chakrabarty et al. 2023). Additionally, many previous works have focused on gradient-based prompt learning methods, such as  (Zhou et al. 2022; Wen et al. 2023). However, these gradient-based methods are not human-readable, which may hinder their application in human-AI collaboration.
@@
Our goal is to generate fine-grained prompts $t_{f}$ from the coarse-grained prompts $t_{c}$ provided by the user while ensuring that the feature space can be understood by the UNet denoising model $\epsilon_{\theta}$, which is designed to generate fine-grained prompts that are model-preferred prompts. To achieve this, we refer to the $\epsilon_{\theta}$ convolutional input as the spatial input $\gamma$ (e.g., $z_{t}$) since convolutions preserve the spatial structure, and to the attention conditioning input as $\psi$ (e.g., $[\tau,E_{T}(T)]$). We train the text encoder $E_{T}$ by minimizing the loss function defined as follows:
@@
In Fig. 3, we visualize the generation results from various models, including GPT-2 (Radford et al. 2019), FLAN-T5 (Chung et al. 2022), GPT-3.5 (OpenAI 2023), and GPT-4 (OpenAI 2023). First, we rewrite the coarse-grained prompts, setting the max tokens to 20. Then, we generate images using Stable Diffusion-v2.1. Our method is capable of producing more visually appealing images. Furthermore, we note that traditional language models, including GPT-2 and FLAN-T5, struggle to comprehend the format of model-preferred prompts in text-to-image tasks. Even when we provide ChatGPT with a prompts format during generation, it still fails to produce satisfactory results. For example, in the case of “a woman in a blue dress”, GPT-4 modifies the original semantics, leading to a significant deviation in the generated result from the original content. In the majority of cases, GPT-2 and FLAN-T5 can only generate short text, even when we attempt to increase the maximum token count. In summary, these issues originate from language models’ limited grasp of image information in text-to-image tasks and their unfamiliarity with the structure of preferred prompts. Our method effectively addresses these issues.
@@
Following (Li et al. 2023) and (Dinh, Nguyen, and Hua 2022), we quantitatively assess the image quality and aesthetic, using the non-reference metrics. Specifically, we choose NIMA (Talebi and Milanfar 2018), MUSIQ (Ke et al. 2021), DB-CNN (Zhang et al. 2020), and TReS (Golestaneh, Dadsetan, and Kitani 2022). We use NIMA-TID, MUSIQ-KonIQ, DB-CNN, and TReS for image quality assessment, while NIMA-AVA and MUSIQ-AVA are used for image aesthetic assessment. NIMA-AVA and MUSIQ-AVA are trained on AVA (Murray, Marchesotti, and Perronnin 2012), MUSIQ-KonIQ and DB-CNN are trained on KonIQ-10K (Hosu et al. 2020), and NIMA-TID is trained on TID2013 (Ponomarenko et al. 2013), following Py-IQA (Chen and Mo 2022).

### 25 Mar 2026 v2

@@ Introduction
Generative foundation models, including language models and text-to-image models, can be prompted to follow user instructions. Recent advancements in text-to-image synthesis such as Stable Diffusion (SD) (Rombach et al. 2022) and Midjourney (Holz 2023) have facilitated the generation of high-fidelity images based on text prompts. Concurrently, recent studies have revealed that prompt design plays a crucial role in determining the quality of the generated images (Ko et al. 2023; Liu et al. 2022). Adjusting the prompt to better reflect the user’s intentions can lead to superior results. This issue is particularly pronounced in text-to-image models, as the capacity of their text encoders is relatively limited, such as CLIP text encoder (Radford et al. 2021) in Stable Diffusion. Empirical observations have also shown that common user inputs are often insufficient to produce aesthetically pleasing images using current models (Hao et al. 2023).
@@
In the specific field of text-to-image generative models, prompt engineering research is nascent, which aims to use carefully selected and combined sentences to achieve a specific visual style in synthesized images (Oppenlaender 2023). The input texts, known as the prompts, direct the model to generate specific images. Manual prompt engineering is a natural method for optimizing prompts. Pavlichenko et al. (2023); Oppenlaender et al. (2023); Liu et al. (2022) explore how to find model-preferred prompts manually. While manual prompt engineering can lead to significant progress, the process of designing prompts requires time and experience, and may not always yield optimal results. Therefore, various methods have focused on automatically searching for prompts through mining (Jiang et al. 2020), parsing (Haviv et al. 2021), text generation (Hao et al. 2023) and LLMs (Zhu et al. 2023; Chakrabarty et al. 2023). Additionally, many previous works have focused on gradient-based prompt learning methods, such as  (Zhou et al. 2022; Wen et al. 2023). However, these gradient-based methods are not human-readable, which may hinder their application in human-AI collaboration.
@@
Our goal is to generate fine-grained prompts $t_{f}$ from the coarse-grained prompts $t_{c}$ provided by the user while ensuring that the feature space can be understood by the UNet denoising model $\epsilon_{\theta}$, which is designed to generate fine-grained prompts that are model-preferred prompts. To achieve this, we refer to the $\epsilon_{\theta}$ convolutional input as the spatial input $\gamma$ (e.g., $z_{t}$) since convolutions preserve the spatial structure, and to the attention conditioning input as $\psi$ (e.g., $\tau,E_{T}(T)]$). We train the text encoder $E_{T}$ by minimizing the loss function defined as follows:
@@
In Fig. 3, we visualize the generation results from various models, including GPT-2 (Radford et al. 2019), FLAN-T5 (Chung et al. 2022), GPT-3.5 (OpenAI 2023), and GPT-4 (OpenAI 2023). First, we rewrite the coarse-grained prompts, setting the max tokens to 20. Then, we generate images using Stable Diffusion-v2.1. Our method is capable of producing visually appealing images. Furthermore, we note that traditional language models, including GPT-2 and FLAN-T5, struggle to comprehend the format of model-preferred prompts in text-to-image tasks. Even when we provide ChatGPT with a prompts format during generation it still fails to produce satisfactory results (bluelovers 2022). For example, in the case of “a woman in a blue dress”, GPT-4 modifies the original semantics, leading to a significant deviation in the generated result from the original content. In most cases, GPT-2 and FLAN-T5 can only generate short text, even when we attempt to increase the maximum token count. In summary, these issues originate from language models’ limited grasp of image information in text-to-image tasks and their unfamiliarity with the structure of preferred prompts. Our method effectively addresses these issues.
@@
Following (Li et al. 2023) and (Dinh et al. 2022), we quantitatively assess the image quality and aesthetic, using the non-reference metrics. Specifically, we choose NIMA (Talebi and Milanfar 2018), MUSIQ (Ke et al. 2021), DB-CNN (Zhang et al. 2020), and TReS (Golestaneh et al. 2022). We use NIMA-TID, MUSIQ-KonIQ, DB-CNN, and TReS for image quality assessment, while NIMA-AVA and MUSIQ-AVA are used for image aesthetic assessment. NIMA-AVA and MUSIQ-AVA are trained on AVA (Murray et al. 2012), MUSIQ-KonIQ and DB-CNN are trained on KonIQ-10K (Hosu et al. 2020), and NIMA-TID is trained on TID2013 (Ponomarenko et al. 2013), following Py-IQA (Chen and Mo 2022).
@@ Acknowledgments
Appendix A: CFP Dataset details
We construct Coarse-Fine Granularity Prompts dataset (CFP) based on Lexica.art (Santana 2022), and organize the data according to DiffusionDB (Wang et al. 2022), as illustrated in Fig. 8. To ensure that the fine-grained prompts in our dataset align with the distribution of training data, we analyze the distribution of fine-grained prompt word lengths, as shown in Fig. 7.
Finally, we obtain a total of 81,910 data instances, each consisting of one fine-grained prompt, one fine-grained image, and three coarse-grained prompts. We split 73,718 data pairs as the training set and 8,192 data instances as the testing set. We also employ an NSFW detector (michellejieli 2022) to filter out fine-grained prompts that contain NSFW elements. We suggest using only the data with scores below 0.9, which amounts to a total of 79,447 data instances. In Table 1 of the main text, we quantitatively analyze the fine-grained prompts and find that they are superior to coarse-grained prompts in terms of both quality and aesthetics. This further demonstrates the necessity of our dataset, and the fine-grained prompts are preferred by models.
Appendix B: Additional Visualizations
We present additional visualization results as shown in Fig. 10 and Fig. 11. As mentioned in Sec. Applications, we propose two strategies for prompt generation. Figure 10 illustrates the case of continuous generation with a maximum token setting of 6. Our UF-FGTG is user-friendly and consistently generates better prompts than other methods. Figure 11 presents additional visual results for Sec. Qualitative Comparison, where the maximum token setting is 20.
Overall, it can be observed that our UF-FGTG generates prompts that correspond to better images than other methods. Our method achieves superior performance in generating images of various categories, including landscapes, animals, abstract scenes, and human portraits.
Figure 7: Word length distribution in fine-grained prompts of the CFP dataset.
Figure 8: The CFP dataset comprises 81,910 data instances, each containing a fine-grained image, a fine-grained prompt, and three coarse-grained prompts of different lengths (1-5 tokens, 6-10 tokens, 11-15 tokens), as well as all model hyperparameters: ‘seed’, ‘step’, ‘CFG scale’, ‘sampler’, and ‘image size’. Additionally, each image has a unique filename. To assist researchers in identifying potentially unsafe or harmful content, we employ state-of-the-art models to calculate an NSFW score for each fine-grained prompt.
Figure 9: Transferability of UF-FGTG across different text-to-image models.
Appendix C: ChatGPT Description
The ChatGPT description, also referred to as “ChatGPT prompt”, is a phrase or instruction given to the ChatGPT AI model to generate a response, which can enhance productivity and creativity. Following (bluelovers 2022), we provide the following description for GPT-3.5 and GPT-4:
Stable Diffusion is an AI art generation model similar to DALLE-2.
Here are some prompts for generating art with Stable Diffusion.
Example:
  • A ghostly apparition drifting through a haunted mansion’s grand ballroom, illuminated by flickering candlelight. Eerie, ethereal, moody lighting.
  • portait of a homer simpson archer shooting arrow at forest monster, front game card, drark, marvel comics, dark, smooth
  • Please see more in (bluelovers 2022).
The prompt should adhere to and include all of the following rules:
  • Prompt should always be written in English, regardless of the input language. Please provide the prompts in English.
  • Each prompt should consist of a description of the scene followed by modifiers divided by commas.
  • When generating descriptions, focus on portraying the visual elements rather than delving into abstract psychological and emotional aspects. Provide clear and concise details that vividly depict the scene and its composition, capturing the tangible elements that make up the setting.
  • The modifiers should alter the mood, style, lighting, and other aspects of the scene.
  • Multiple modifiers can be used to provide more specific details.
  • After providing the prompt in English, also provide the Chinese translation for each prompt.
I want you to write me each prompt exactly about a list of 1 prompt following the rule, and please give me the final 1 prompt.
Appendix D: Transferability of UF-FGTG
Fig. 9 shows our approach’s generalizability across various text-to-image models. Our model, initially trained on Stable Diffusion-v2.1, leverages its inherent robustness to generate effective prompts. We hypothesize this robustness allows our model to produce high-quality images even when applied to other models. To verify this, we test our model on Midjourney (Holz 2023), a text-to-image model distinct from Stable Diffusion. Our results confirm that our model generates prompts leading to high-quality images on Midjourney. This suggests our model’s prompt-generation ability extends beyond Stable Diffusion to other models.
Overall, our model’s generated prompts can also produce high-quality images on other text-to-image models. This indicates that our method has learned a generalizable representation of text-to-image prompts and can be widely applied to help users generate model-preferred prompts
Figure 10: Comparison of continuous generation prompts generated by FLAN-T5, GPT-2, GPT3.5, GPT-4, and our UF-FGTG, with corresponding images generated by Stable Diffusion-v2.1 (max token set to 6). Our UF-FGTG generates more reasonable and user-friendly prompts at each step, resulting in visually superior images compared to other methods. Origin prompt: “cute cat”.
Figure 11: Comparison of prompts generated by FLAN-T5, GPT-2, GPT3.5, GPT-4, and our UF-FGTG, with corresponding images generated by Stable Diffusion-v2.1 (max token set to 20). Our UF-FGTG outperforms other methods significantly in various categories, including animals, humans, landscapes, and plants.